OpenStudy (anonymous):

$1400 is deposited in an account with an interest rate of r% per year, compounded monthly. At the end of 8 years, the balance in the account is given by A = 1400(1 +r/1200)^96. Find the rate of change of A with respect to r when r = 8.

OpenStudy (rob1525):

Im wondering if their asking to calculate the first derivative of the function then pugging in 8 for the derived function. Does that sound familiar?

OpenStudy (anonymous):

im not sure, they're asking for rate of change of A, i think chain rule is involved wihtin this problem so it would be derived?

OpenStudy (anonymous):

This question seems relatively simple if you are know how to use derivatives. Just take the derivative with respect to r for that equation \[\frac{ dA }{ dr } = 1400*96*(1+\frac{ r }{ 1200 })^{95}*\frac{ 1 }{ 1200 }\] then simplify

OpenStudy (anonymous):

i forgot the 1/1200 at the end, simple formula mistakes... thank you

OpenStudy (anonymous):

your welcome do not forget to close the question and medal.

OpenStudy (anonymous):

forget the medal part was working on my own calculus question